# A Comparison Between Invariant and Equivariant Classical and Quantum Graph Neural Networks

### v2

@@ Abstract
Machine learning algorithms are heavily relied on to understand the vast amounts of data from high-energy particle collisions at the CERN Large Hadron Collider (LHC). The data from such collision events can naturally be represented with graph structures. Therefore, deep geometric methods, such as graph neural networks (GNNs), have been leveraged for various data analysis tasks in high-energy physics. One typical task is jet tagging, where jets are viewed as point clouds with distinct features and edge connections between their constituent particles. The increasing size and complexity of the LHC particle datasets, as well as the computational models used for their analysis, greatly motivate the development of alternative fast and efficient computational paradigms such as quantum computation. In addition, to enhance the validity and robustness of deep networks, one can leverage the fundamental symmetries present in the data through the use of invariant inputs and equivariant layers. In this paper, we perform a fair and comprehensive comparison between classical graph neural networks (GNNs) and equivariant graph neural networks (EGNNs) and their quantum counterparts: quantum graph neural networks (QGNNs) and equivariant quantum graph neural networks (EQGNN). The four architectures were benchmarked on a binary classification task to classify the parton-level particle initiating the jet. Based on their AUC scores, the quantum networks were shown to outperform the classical networks. However, seeing the computational advantage of the quantum networks in practice may have to wait for the further development of quantum technology and its associated APIs.
@@ 1 Introduction
Through the measurement of byproducts of particle collisions, the Large Hadron Collider (LHC) collects a substantial amount of information about the fundamental particles and their interactions. The data produced from these collisions can be analyzed using various supervised and unsupervised machine learning methods Andreassen et al. (2019); Shlomi et al. (2020); Mikuni and Canelli (2021); Mokhtar et al. (2022); Mikuni and Canelli (2020). Jet tagging is a key task in high energy physics which seeks to identify the likely parton-level particle from which the jet originated. By viewing individual jets as point clouds with distinct features and edge connections between the constituent particles which make them up, a graph neural network (GNN) has been considered as a well-suited architecture for jet tagging Mikuni and Canelli (2021); Shlomi et al. (2020).
@@
Classified as deep geometric networks, GNNs have the potential to draw inferences about a graph structure, including the interactions among elements in the graph Veličković (2023); Zhou et al. (2020). Graph neural networks are typically thought of as generalizations of convolutional neural networks (CNNs), which are predominantly used for image recognition, pattern recognition, and computer vision Kipf and Welling (2016, 2017). This can be attributed to the fact that, in an image, each pixel is connected to its nearest neighboring pixels, whereas in a general dataset, one would ideally like to construct an arbitrary graph structure among the samples. Many instances in nature can be described well in terms of graphs, including molecules, maps, social networks, and the brain. For example, in molecules, the nodal data can be attributed to the atoms, the edges can be characterized as the strength of the bond between atoms, and the features embedded within each node can be the atom characteristics, such as the reactivity.
@@
Generally, graphically-structured problems involve unordered sets of elements with learnable embedding of the input features. Useful information can be extracted from such graph structured data by the embedding it within GNNs. Many subsequent developments have been made to progress GNNs since their first implementation in 2005, including graph convolutional, recurrent, message passing, graph attention, and graph transformer, architectures Veličković (2023); Shlomi et al. (2020); Gilmer et al. (2017); Veličković et al. (2018).
@@
To enhance the validity and robustness of deep networks, invariant and equivariant networks have been constructed to learn the symmetries embedded within a dataset by preserving an oracle in the former and by enforcing weight sharing across filter orientations in the latter Lim and Nelson (2022); Ecker et al. (2019). Utilizing analytical invariant quantities characteristic of physical symmetry representations, computational methods have successfully rediscovered fundamental Lie group structures, such as the $SO(n)$, $SO(1,3)$, and $U(n)$ groups Forestano et al. (2023b, d, c, a). Non-linear symmetry discovery methods have also been applied across classification tasks in data domains Roman et al. (2023). The simplest and most useful embedded symmetry transformations include translations, rotations, and reflections, which have been the primary focus in invariant (IGNN) and equivariant graph neural networks (EGNN) Maron et al. (2019); Satorras et al. (2021).
@@
The learned representations from the collection of these network components can be used to understand unobservable causal factors, uncover fundamental physical principles governing these processes, and possibly even discover statistically significant hidden anomalies. However, with increasing amounts of available data and the computational cost of these deep learning networks, large computing resources will be required to efficiently run these machine learning algorithms. The extension of classical networks, which relies on bit-wise computation, to quantum networks, which relies on qubit-wise computation, is already underway as a solution to this complexity problem. Many API’s, including Xanadu’s Pennylane, Google’s Cirq, and IBM’s Qiskit, have been developed to allow for the testing of quantum circuits and quantum machine learning algorithms running on these quantum devices. In the quantum graph structure, classical nodes can be mapped to quantum states, real valued features to the complex valued characteristics of the states, edges to the interactions between states, and edge attributes to the strength of the interactions between the quantum states. Through a well-defined Hamiltonian operator, the larger structure of a classical model can then be embedded into the quantum model. This parameterized Hamiltonian determines the evolution of the quantum system by acting on the quantum states within the graph. Following several layers of application, a final state measurement of the quantum system can then be made to reach a final prediction. The theory and application of unsupervised and supervised learning tasks involving quantum graph neural networks (QGNNs), quantum graph recurrent neural networks (QGRNNs), and quantum graph convolutional neural networks (QGCNNs) have already been developed Beer et al. (2023); Verdon et al. (2019). Improvements of these models to arbitrarily sized graphs have been made with the implemenation of ego-graph based quantum graph neural networks (egoQGNN) Ai et al. (2022). Quantum analogs of other advanced classical architectures, including generative adversarial networks (GANs), transformers, natural language processors (NLPs), and equivariant networks, have also been proposed Beer et al. (2023); Niu et al. (2022); Chu et al. (2023); Sipio et al. (2021); Cherrat et al. (2023); Meyer et al. (2023); Nguyen et al. (2022); Schatzki et al. (2022).
@@
With the rapid development of quantum deep learning, this work intends to offer a fair and comprehensive comparison between GNNs and their quantum counterparts. To classify whether a particle jet has originated from a quark or a gluon, a binary classification task was carried out across four different architectures. These architectures include a GNN, SE(2) EGNN, QGNN, and permutation EQGNN. Each quantum model was fine tuned to have an analogous structure to its classical form. In order to provide a fair comparison, all models used similar hyperparameters, as well as a similar number of total trainable parameters. The final results across each architecture using identical training, validation, and testing sets were recorded. We find that the QGNN and EQGNN outperform their classical analogs on the particular binary classification task described above. Although these results seem promising for the future of quantum computing, further development of quantum application programming interfaces (APIs) is required to allow for more general implementations of quantum architectures.
@@ 2 Data
The jet-tagging binary classification task is illustrated with the high energy physics (HEP) dataset Pythia8 Quark and Gluon Jets for Energy Flow Komiske et al. (2019), which contains two million jets split equally into one million quark jets and one million gluon jets. These jets resulted from LHC collisions with total center of mass energy $\sqrt{s}=14$ TeV and were selected to have transverse momenta $p_{T}^{jet}$ between $500$ to $550$ GeV and rapidities $|y^{jet}|<1.7$. The jet kinematic distributions are shown in Fig. 1. For each jet $\alpha$, the classification label is provided as either a quark with $y_{\alpha}=1$ or a gluon with $y_{\alpha}=0$, and each particle $i$ within the jet is listed with its transverse momentum $p^{(i)}_{T,\alpha}$, rapidity $y_{\alpha}^{(i)}$, azimuthal angle $\phi_{\alpha}^{(i)}$, and PDG id $I_{\alpha}^{(i)}$.
@@
2.1 Graphically structured data
A graph $\mathcal{G}$ is typically defined as a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$, i.e. $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. Each node $v^{(i)}\in\mathcal{V}$ is connected to its neighboring nodes $v^{(j)}$ via edges $e^{(ij)}\in\mathcal{E}$. In our case, each jet $\alpha$ can be considered as a graph $\mathcal{J}_{\alpha}$ composed of the jet’s constituent particles as nodes $v_{\alpha}^{(i)}$ with node features $h_{\alpha}^{(i)}$ and edge connections $e_{\alpha}^{(ij)}$ between the nodes in $\mathcal{J}_{\alpha}$ with edge features $a_{\alpha}^{(ij)}$. It should be noted that the number of nodes within a graph can vary. This is especially true for the case of particle jets, where the number of particles within each jet can vary greatly. Each jet $\mathcal{J}_{\alpha}$ can be considered as a collection of $m_{\alpha}$ particles with $l$ distinct features per particle. An illustration of graphically structured data and an example jet in the $(\phi,y)$ plane can be seen in Fig. 2.
@@ 2.2 Feature engineering
We use the Particle package Rodrigues and Schreiner (2023) to find the particle masses $m_{\alpha}^{(i)}$ from the respective particle id’s $I_{\alpha}^{(i)}$. From the available kinematic information for each particle $i$, we construct new physically meaningful kinematic variables Franceschini et al. (2022), which will be used as additional features in the analysis. In particular, we consider the transverse mass $m_{T,\alpha}^{(i)}$, the energy $E_{\alpha}^{(i)}$, and the Cartesian momentum components, $p_{x,\alpha}^{(i)}$, $p_{y,\alpha}^{(i)}$, and $p_{z,\alpha}^{(i)}$, defined respectively as
$m_{T,\alpha}^{(i)}$ $=\sqrt{m_{\alpha}^{(i)2}+p_{T,\alpha}^{(i)2}},\qquad E_{\alpha}^{(i)}=m_{T,\alpha}^{(i)}\text{cosh}(y_{\alpha}^{(i)}),$ (1a)
$p_{x,\alpha}^{(i)}$ $=p_{T,\alpha}^{(i)}\text{cos}(\phi_{\alpha}^{(i)}),\qquad p_{y,\alpha}^{(i)}=p_{T,\alpha}^{(i)}\text{sin}(\phi_{\alpha}^{(i)}),\qquad p_{z,\alpha}^{(i)}=m_{T,ij}\text{sinh}(y_{\alpha}^{(i)}).$ (1b)
The original kinematic information in the dataset is then combined with the additional kinematic variables (2.2) into a feature set $h_{\alpha}^{(il)}$, $l=0,1,2,\ldots,7$, as follows
@@
These features were then max scaled by their maximum value across all jets $\alpha$ and particles $i$, i.e. $h_{\alpha}^{(il)}\to h_{\alpha}^{(il)}/max_{\alpha,i}(h_{\alpha}^{(il)})$.
@@ 2.3 Training, Validation, and Testing Sets
We consider jets with at least $10$ particles. This leaves us with $N=1,997,445$ jets, $997,805$ of which are quark jets. While the classical GNN is more flexible in terms of its hidden features, the size of the quantum state and the Hamiltonian scale as $2^{n}$, where $n$ is the number of qubits. As we shall see, the number of qubits will be given by the number of nodes $n_{\alpha}$ in the graph, i.e. the number of particles in the jet. Therefore, jets with large particle multiplicity will require prohibitively complex quantum networks. This is why, we limit ourselves to the case of $n_{\alpha}=3$, and for each jet, only consider the three highest $p_{T}$ particles. In other words, each graph contains the set $\mathbf{h}_{\alpha}=(\mathbf{h}_{\alpha}^{(1)},\mathbf{h}_{\alpha}^{(2)},\mathbf{h}_{\alpha}^{(3)})$ where each $\mathbf{h}_{\alpha}^{(i)}\in\mathbb{R}^{8}$ and $\mathbf{h}_{\alpha}\in\mathbb{R}^{3\times 8}$. For training, we randomly picked $N=12,500$ jets and used the first $10,000$ for training, the next $1,250$ for validation, and the last $1,250$ for testing. These sets happened to contain $4982$, $658$, and $583$ quark jets, respectively.
@@ 3 Models
The four different models described below, including a GNN, EGNN, QGNN, and EQGNN, were constructed to perform graph classification. The binary classification task was to determine whether a jet $\mathcal{J}_{\alpha}$ originated from a quark or a gluon.
@@ 3.1 Invariance and Equivariance
By making a network invariant or equivariant to particular symmetries within a dataset, a more robust architecture can be developed. In order to introduce invariance and equivariance, one must assume or learn a certain symmetry group $G$ of transformations on the dataset. A function $\varphi:X\to Y$ is equivariant with respect to a set of group transformations $T_{g}:X\to X$, $g\in G$, acting on the input vector space $X$, if there exists a set of transformations $S_{g}:Y\to Y$ which similarly transform the output space $Y$, i.e.
@@
A model is said to be invariant when for all $g\in G$, $S_{g}$ becomes the set containing only the trivial mapping, i.e. $S_{g}=\{\mathbb{I}_{G}\}$, where $\mathbb{I}_{G}\in G$ is the identity element of the group $G$ Lim and Nelson (2022); Esteves (2020).
Invariance has been shown to perform better as an input embedding, whereas equivariance can be more easily incorporated into the model layers. For each model, the invariant component corresponds to the translational and rotational invariant embedding distance $\varphi\equiv\Delta R_{\alpha}^{(ij)}$, where in this case $\varphi:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}$, which makes up the edge attribute matrix $a_{\alpha}^{(ij)}$, as defined in Eq. 3. This distance is used as opposed to solely incorporating the raw coordinates. Equivariance has been accomplished through the use of simple nontrivial functions along with higher order methods involving the use of spherical harmonics to embed the equivariance within the network Murnane et al. (2023); Worrall et al. (2017). Equivariance takes different forms in each model presented here.

### v3

# quantum computing; deep learning; quantum machine learning; equivariance; invariance; supervised learning; classification; particle physics; Large Hadron Collider

@@ Abstract
Machine learning algorithms are heavily relied on to understand the vast amounts of data from high-energy particle collisions at the CERN Large Hadron Collider (LHC). The data from such collision events can naturally be represented with graph structures. Therefore, deep geometric methods, such as graph neural networks (GNNs), have been leveraged for various data analysis tasks in high-energy physics. One typical task is jet tagging, where jets are viewed as point clouds with distinct features and edge connections between their constituent particles. The increasing size and complexity of the LHC particle datasets, as well as the computational models used for their analysis, have greatly motivated the development of alternative fast and efficient computational paradigms such as quantum computation. In addition, to enhance the validity and robustness of deep networks, we can leverage the fundamental symmetries present in the data through the use of invariant inputs and equivariant layers. In this paper, we provide a fair and comprehensive comparison of classical graph neural networks (GNNs) and equivariant graph neural networks (EGNNs) and their quantum counterparts: quantum graph neural networks (QGNNs) and equivariant quantum graph neural networks (EQGNN). The four architectures were benchmarked on a binary classification task to classify the parton-level particle initiating the jet. Based on their area under the curve (AUC) scores, the quantum networks were found to outperform the classical networks. However, seeing the computational advantage of quantum networks in practice may have to wait for the further development of quantum technology and its associated application programming interfaces (APIs).
keywords:
quantum computing; deep learning; quantum machine learning; equivariance; invariance; supervised learning; classification; particle physics; Large Hadron Collider
[…]
@@ 1 Introduction
Through the measurement of the byproducts of particle collisions, the Large Hadron Collider (LHC) collects a substantial amount of information about fundamental particles and their interactions. The data produced from these collisions can be analyzed using various supervised and unsupervised machine learning methods Andreassen et al. (2019); Shlomi et al. (2020); Mikuni and Canelli (2021); Mokhtar et al. (2022); Mikuni and Canelli (2020). Jet tagging is a key task in high-energy physics, which seeks to identify the likely parton-level particle from which the jet originated. By viewing individual jets as point clouds with distinct features and edge connections between their constituent particles, a graph neural network (GNN) is considered a well-suited architecture for jet tagging Mikuni and Canelli (2021); Shlomi et al. (2020).
@@
Classified as deep geometric networks, GNNs have the potential to draw inferences about a graph structure, including the interactions among the elements in the graph Veličković (2023); Zhou et al. (2020). Graph neural networks are typically thought of as generalizations of convolutional neural networks (CNNs), which are predominantly used for image recognition, pattern recognition, and computer vision Kipf and Welling (2016, 2017). This can be attributed to the fact that in an image, each pixel is connected to its nearest neighboring pixels, whereas in a general dataset, one would ideally like to construct an arbitrary graph structure among the samples. Many instances in nature can be described well in terms of graphs, including molecules, maps, social networks, and the brain. For example, in molecules, the nodal data can be attributed to the atoms, the edges can be characterized as the strength of the bond between atoms, and the features embedded within each node can be the atom’s characteristics, such as reactivity.
@@
Generally, graphically structured problems involve unordered sets of elements with a learnable embedding of the input features. Useful information can be extracted from such graphically structured data by embedding them within GNNs. Many subsequent developments have been made to GNNs since their first implementation in 2005. These developments have included graph convolutional, recurrent, message passing, graph attention, and graph transformer architectures Veličković (2023); Shlomi et al. (2020); Gilmer et al. (2017); Veličković et al. (2018).
@@
To enhance the validity and robustness of deep networks, invariant and equivariant networks have been constructed to learn the symmetries embedded within a dataset by preserving an oracle in the former and by enforcing weight sharing across filter orientations in the latter Lim and Nelson (2022); Ecker et al. (2019). Utilizing analytical invariant quantities characteristic of physical symmetry representations, computational methods have successfully rediscovered fundamental Lie group structures, such as the $SO(n)$, $SO(1,3)$, and $U(n)$ groups Forestano et al. (2023a, b, c, d). Nonlinear symmetry discovery methods have also been applied to classification tasks in data domains Roman et al. (2023). The simplest and most useful embedded symmetry transformations include translations, rotations, and reflections, which have been the primary focus in invariant (IGNN) and equivariant (EGNN) graph neural networks Maron et al. (2019); Gong et al. (2022); Satorras et al. (2021).
@@
The learned representations from the collection of these network components can be used to understand unobservable causal factors, uncover fundamental physical principles governing these processes, and possibly even discover statistically significant hidden anomalies. However, with increasing amounts of available data and the computational cost of these deep learning networks, large computing resources will be required to efficiently run these machine learning algorithms. The extension of classical networks, which rely on bit-wise computation, to quantum networks, which rely on qubit-wise computation, is already underway as a solution to this complexity problem. Due to superposition and entanglement among qubits, quantum networks are able to store the equivalent of $2^{n}$ characteristics from $n$ two-dimensional complex vectors. In other words, while the expressivity of the classical network scales linearly, that of the quantum network scales exponentially with the sample size $n$ Preskill (2018). Many APIs, including Xanadu’s Pennylane, Google’s Cirq, and IBM’s Qiskit, have been developed to allow for the testing of the quantum circuits and quantum machine learning algorithms running on these quantum devices.
In the quantum graph structure, classical nodes can be mapped to the quantum states of the qubits, real-valued features to the complex-valued entries of the states, edges to the interactions between states, and edge attributes to the strength of the interactions between the quantum states. Through a well-defined Hamiltonian operator, the larger structure of a classical model can then be embedded into the quantum model. The unitary operator constructed from this parameterized Hamiltonian determines the temporal evolution of the quantum system by acting on the fully entangled quantum state of the graph. Following several layers of application, a final state measurement of the quantum system can then be made to reach a final prediction. The theory and application of unsupervised and supervised learning tasks involving quantum graph neural networks (QGNNs), quantum graph recurrent neural networks (QGRNNs), and quantum graph convolutional neural networks (QGCNNs) have already been developed Beer et al. (2023); Verdon et al. (2019). Improvements to these models to arbitrarily sized graphs have been made with the implementation of ego-graph-based quantum graph neural networks (egoQGNNs) Ai et al. (2022). Quantum analogs of other advanced classical architectures, including generative adversarial networks (GANs), transformers, natural language processors (NLPs), and equivariant networks, have also been proposed Beer et al. (2023); Niu et al. (2022); Chu et al. (2023); Sipio et al. (2021); Cherrat et al. (2023); Meyer et al. (2023); Nguyen et al. (2022); Schatzki et al. (2022).
@@
With the rapid development of quantum deep learning, this paper intends to offer a fair and comprehensive comparison between classical GNNs and their quantum counterparts. To classify whether a particle jet has originated from a quark or a gluon, a binary classification task was carried out using four different architectures. These architectures included a GNN, SE(2) EGNN, QGNN, and permutation EQGNN. Each quantum model was fine tuned to have an analogous structure to its classical form. In order to provide a fair comparison, all models used similar hyperparameters as well as a similar number of total trainable parameters. The final results across each architecture were recorded using identical training, validation, and testing sets. We found that QGNN and EQGNN outperformed their classical analogs on the particular binary classification task described above. Although these results seem promising for the future of quantum computing, the further development of quantum APIs is required to allow for more general implementations of quantum architectures.
@@ 2 Data
The jet tagging binary classification task is illustrated with the high-energy physics (HEP) dataset Pythia8 Quark and Gluon Jets for Energy Flow Komiske et al. (2019). This dataset contains data from two million particle collision jets split equally into one million jets that originated from a quark and one million jets that originated from a gluon. These jets resulted from LHC collisions with total center of mass energy $\sqrt{s}=14$ TeV and were selected to have transverse momenta $p_{T}^{jet}$ between $500$ to $550$ GeV and rapidities $|y^{jet}|<1.7$. The jet kinematic distributions are shown in Figure 1. For each jet $\alpha$, the classification label is provided as either a quark with $y_{\alpha}=1$ or a gluon with $y_{\alpha}=0$. Each particle $i$ within the jet is listed with its transverse momentum $p^{(i)}_{T,\alpha}$, rapidity $y_{\alpha}^{(i)}$, azimuthal angle $\phi_{\alpha}^{(i)}$, and PDG id $I_{\alpha}^{(i)}$.
{adjustwidth}
[…]
@@
2.1 Graphically Structured Data
A graph $\mathcal{G}$ is typically defined as a set of nodes $\mathcal{V}$ and edges $\mathcal{E}$, i.e., $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$. Each node $v^{(i)}\in\mathcal{V}$ is connected to its neighboring nodes $v^{(j)}$ via edges $e^{(ij)}\in\mathcal{E}$. In our case, each jet $\alpha$ can be considered as a graph $\mathcal{J}_{\alpha}$ composed of the jet’s constituent particles as the nodes $v_{\alpha}^{(i)}$ with node features $h_{\alpha}^{(il)}$ and edge connections $e_{\alpha}^{(ij)}$ between the nodes in $\mathcal{J}_{\alpha}$ with edge features $a_{\alpha}^{(ij)}$. It should be noted that the number of nodes within a graph can vary. This is especially true for the case of particle jets, where the number of particles within each jet can vary greatly. Each jet $\mathcal{J}_{\alpha}$ can be considered as a collection of $m_{\alpha}$ particles with $l$ distinct features per particle. An illustration of graphically structured data and an example jet in the $(\phi,y)$ plane are shown in Figure 2.
@@ 2.2 Feature Engineering
We used the Particle package Rodrigues and Schreiner (2023) to find the particle masses $m_{\alpha}^{(i)}$ from the respective particle IDs $I_{\alpha}^{(i)}$. From the available kinematic information for each particle $i$, we constructed new physically meaningful kinematic variables Franceschini et al. (2022), which were used as additional features in the analysis. In particular, we considered the transverse mass $m_{T,\alpha}^{(i)}$, the energy $E_{\alpha}^{(i)}$, and the Cartesian momentum components, $p_{x,\alpha}^{(i)}$, $p_{y,\alpha}^{(i)}$, and $p_{z,\alpha}^{(i)}$, defined, respectively, as
$m_{T,\alpha}^{(i)}$ $=\sqrt{m_{\alpha}^{(i)2}+p_{T,\alpha}^{(i)2}},\qquad E_{\alpha}^{(i)}=m_{T,\alpha}^{(i)}\text{cosh}(y_{\alpha}^{(i)}),$ (1)
$p_{x,\alpha}^{(i)}$ $=p_{T,\alpha}^{(i)}\text{cos}(\phi_{\alpha}^{(i)}),\qquad p_{y,\alpha}^{(i)}=p_{T,\alpha}^{(i)}\text{sin}(\phi_{\alpha}^{(i)}),\qquad p_{z,\alpha}^{(i)}=m_{T,ij}\text{sinh}(y_{\alpha}^{(i)}).$
The original kinematic information in the dataset was then combined with the additional kinematic variables (1) into a feature set $h_{\alpha}^{(il)}$, $l=0,1,2,\ldots,7$, as follows:
@@
These features were then max-scaled by their maximum value across all jets $\alpha$ and particles $i$, i.e., $h_{\alpha}^{(il)}\to h_{\alpha}^{(il)}/max_{\alpha,i}(h_{\alpha}^{(il)})$.
@@ 2.3 Training, Validation, and Testing Sets
We considered jets with at least $10$ particles. This left us with $N$ = 1,997,445 jets, 997,805 of which were quark jets. While the classical GNN is more flexible in terms of its hidden features, the size of the quantum state and the Hamiltonian scale as $2^{n}$, where $n$ is the number of qubits. As we shall see, the number of qubits is given by the number of nodes $n_{\alpha}$ in the graph, i.e., the number of particles in the jet. Therefore, jets with large particle multiplicity require prohibitively complex quantum networks. Thus, we limited ourselves to the case of $n_{\alpha}=3$ particles per jet by only considering the three highest momenta $p_{T}$ particles within each jet. In other words, each graph contained the set $\mathbf{h}_{\alpha}=(\mathbf{h}_{\alpha}^{(1)},\mathbf{h}_{\alpha}^{(2)},\mathbf{h}_{\alpha}^{(3)})$, where each $\mathbf{h}_{\alpha}^{(i)}\in\mathbb{R}^{8}$ and $\mathbf{h}_{\alpha}\in\mathbb{R}^{3\times 8}$. For training, we randomly picked $N$ = 12,500 jets and used the first 10,000 for training, the next 1250 for validation, and the last 1250 for testing. These sets happened to contain $4982$, $658$, and $583$ quark jets, respectively.
@@ 3 Models
The four different models described below, including a GNN, an EGNN, a QGNN, and an EQGNN, were constructed to perform graph classification. The binary classification task was determining whether a jet $\mathcal{J}_{\alpha}$ originated from a quark or a gluon.
@@ 3.1 Invariance and Equivariance
By making a network invariant or equivariant to particular symmetries within a dataset, a more robust architecture can be developed. In order to introduce invariance and equivariance, one must assume or learn a certain symmetry group $G$ of transformations on the dataset. A function $\varphi:X\to Y$ is equivariant with respect to a set of group transformations $T_{g}:X\to X$, $g\in G$, acting on the input vector space $X$, if there exists a set of transformations $S_{g}:Y\to Y$ that similarly transform the output space $Y$, i.e.,
@@
A model is said to be invariant when, for all $g\in G$, $S_{g}$ becomes the set containing only the trivial mapping, i.e., $S_{g}=\{\mathbb{I}_{G}\}$, where $\mathbb{I}_{G}\in G$ is the identity element of the group $G$ Lim and Nelson (2022); Esteves (2020).
Invariance performs better as an input embedding, whereas equivariance can be more easily incorporated into the model layers. For each model, the invariant component corresponds to the translational and rotational invariant embedding distance $\varphi\equiv\Delta R_{\alpha}^{(ij)}$. Here, the function $\varphi:\mathbb{R}^{2}\times\mathbb{R}^{2}\to\mathbb{R}$ makes up the edge attribute matrix $a_{\alpha}^{(ij)}$, as defined in Equation (3). This distance is used as opposed to solely incorporating the raw coordinates. Equivariance has been accomplished through the use of simple nontrivial functions along with higher-order methods involving the use of spherical harmonics to embed the equivariance within the network Murnane et al. (2023); Worrall et al. (2017). Equivariance takes different forms in each model presented here.